# Orientation Matters: Making 3D Generative Models Orientation-Aligned

### v1

@@ B.2 Orientation-aligned Object Generation
To investigate whether our orientation alignment fine-tuning affects the performance of the original pre-trained checkpoint, we randomly select 46 objects across 46 categories in the Toys4k [38] dataset. As shown in Table 6, our method not only maintains performance but may even enhance the quality of the pre-trained 3D generative models. This improvement is likely due to the high fidelity of our manually corrected Objaverse-OA dataset.
@@ B.3 Orientation Estimation
| FSDetView [57] | - / - | 40.00 / 58.52 | 11.11 / 114.89 | - / - | 20.00 / 75.71 | 40.00 / 58.07 |
| Orient Anything [51] (Vit-S) | 30.00 / 70.48 | 20.00 / 63.91 | 44.44 / 73.69 | 33.33 / 47.79 | 100.00 / 11.64 | 80.00 / 29.49 |
| Orient Anything [51] (Vit-L) | 80.00 / 39.03 | 40.00 / 74.73 | 44.44 / 87.91 | 88.89 / 15.63 | 80.00 / 20.32 | 80.00 / 18.73 |
@@
| FSDetView [57] | - / - | 10.00 / 96.30 | - / - | - / - | - / - | - / - |
| Orient Anything [51] (Vit-S) | 55.56 / 35.79 | 55.00 / 39.50 | 25.00 / 59.75 | 33.33 / 61.73 | 10.00 / 75.44 | 50.00 / 29.89 |
| Orient Anything [51] (Vit-L) | 77.78 / 19.83 | 90.00 / 15.53 | 50.00 / 34.62 | 50.00 / 40.90 | 50.00 / 42.89 | 100.00 / 17.54 |
@@
| FSDetView [57] | - / - | - / - | - / - | - / - | - / - | - / - |
| Orient Anything [51] (Vit-S) | 50.00 / 32.32 | 22.22 / 53.95 | 30.00 / 86.84 | 20.00 / 69.35 | 50.00 / 45.98 | 20.00 / 59.59 |
| Orient Anything [51] (Vit-L) | 70.00 / 22.11 | 88.89 / 16.71 | 30.00 / 50.37 | 50.00 / 48.06 | 70.00 / 28.76 | 40.00 / 47.36 |
@@
| FSDetView [57] | - / - | - / - | - / - | 0.00 / 110.48 | - / - | 28.57 / 49.01 |
| Orient Anything [51] (Vit-S) | 77.78 / 31.87 | 50.00 / 48.89 | 0.00 / 51.62 | 30.00 / 67.83 | 0.00 / 101.22 | 14.29 / 53.71 |
| Orient Anything [51] (Vit-L) | 66.67 / 22.15 | 90.00 / 21.53 | 80.00 / 24.69 | 50.00 / 35.85 | 50.00 / 74.92 | 14.29 / 54.06 |
@@
| FSDetView [57] | - / - | 62.50 / 42.29 | - / - | - / - | - / - | - / - |
| Orient Anything [51] (Vit-S) | 20.00 / 73.80 | 75.00 / 19.83 | 50.00 / 36.90 | 10.00 / 56.94 | 90.00 / 19.91 | 25.00 / 101.14 |
| Orient Anything [51] (Vit-L) | 80.00 / 36.15 | 50.00 / 70.09 | 70.00 / 28.20 | 40.00 / 38.71 | 70.00 / 21.11 | 50.00 / 51.71 |
@@
| FSDetView [57] | 10.00 / 80.55 | - / - | - / - | - / - | 33.33 / 102.13 | - / - |
| Orient Anything [51] (Vit-S) | 20.00 / 43.97 | 57.14 / 30.07 | 0.00 / 116.94 | 50.00 / 49.04 | 50.00 / 64.67 | 30.00 / 50.62 |
| Orient Anything [51] (Vit-L) | 40.00 / 29.71 | 71.43 / 14.63 | 0.00 / 89.58 | 56.25 / 33.90 | 50.00 / 61.28 | 80.00 / 30.68 |
@@
| FSDetView [57] | - / - | - / - | - / - | - / - | 40.00 / 98.08 | 15.00 / 107.20 |
| Orient Anything [51] (Vit-S) | 50.00 / 66.47 | 62.50 / 63.46 | 20.00 / 50.26 | 75.00 / 34.54 | 50.00 / 62.30 | 75.00 / 33.12 |
| Orient Anything [51] (Vit-L) | 50.00 / 23.66 | 87.50 / 32.48 | 40.00 / 45.30 | 100.00 / 16.06 | 50.00 / 32.59 | 65.00 / 39.84 |
@@
| FSDetView [57] | - / - | 0.00 / 93.15 | - / - | - / - | - / - | |
| Orient Anything [51] (Vit-S) | 75.00 / 32.67 | 0.00 / 117.58 | 80.00 / 23.95 | 0.00 / 95.34 | 44.44 / 47.62 | |
| Orient Anything [51] (Vit-L) | 75.00 / 37.13 | 50.00 / 70.34 | 100.00 / 11.94 | 30.00 / 45.89 | 44.44 / 39.14 | |
@@
| FSDetView [57] | 9.09 / 90.67 | 4.76 / 81.40 | 0.00 / 81.72 | 54.84 / 40.24 | 0.00 / 93.31 |
| Orient Anything [51] (Vit-S) | 4.55 / 83.23 | 0.00 / 78.44 | 6.67 / 74.79 | 3.23 / 87.56 | 3.03 / 74.11 |
| Orient Anything [51] (Vit-L) | 9.09 / 78.51 | 0.00 / 78.95 | 13.33 / 76.79 | 6.45 / 81.73 | 3.03 / 85.85 |
@@
| FSDetView [57] | 0.00 / 96.13 | 0.00 / 103.08 | |||
| Orient Anything [51] (Vit-S) | 0.00 / 90.53 | 2.63 / 76.98 | |||
| Orient Anything [51] (Vit-L) | 17.24 / 70.88 | 15.79 / 68.49 | |||
@@ B.4 Impact of Occlusion
As illustrated in Figure 22, our object generation module is capable of producing reasonable results under partial occlusion, with minimal performance degradation. However, in scenarios involving severe occlusion, we recommend using the method proposed in [55], which is specifically designed for occlusion-robustness, as the pre-trained checkpoint prior to applying our orientation-alignment fine-tuning. It is also important to note that our orientation estimation approach has not been designed to handle occluded inputs. Further research is necessary to extend its applicability to occlusion-prone scenarios.
@@ Appendix C License and Border Impact
| Objaverse [6] | ODC-By v1.0 license | texthttps://objaverse.allenai.org/ |
| Trellis [56] | MIT License | texthttps://github.com/microsoft/TRELLIS |
| Wonder3D [22] | MIT License | texthttps://github.com/xxlong0/Wonder3D/tree/main |
| FoundationPose [53] | NVIDIA Source Code License | texthttps://github.com/NVlabs/FoundationPose |
| LGM [43] | MIT License | texthttps://github.com/3DTopia/LGM |

### v2

@@ B.2 Orientation-aligned Object Generation
To investigate whether our orientation alignment fine-tuning affects the performance of the original pre-trained checkpoint, we randomly select 46 objects across 46 categories in the Toys4k [Stojanov2021UsingST] dataset. As shown in Table 6, our method not only maintains performance but may even enhance the quality of the pre-trained 3D generative models. This improvement is likely due to the high fidelity of our manually corrected Objaverse-OA dataset.
@@ B.3 Orientation Estimation
| FSDetView [Xiao2020FewShotOD] | - / - | 40.00 / 58.52 | 11.11 / 114.89 | - / - | 20.00 / 75.71 | 40.00 / 58.07 |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 30.00 / 70.48 | 20.00 / 63.91 | 44.44 / 73.69 | 33.33 / 47.79 | 100.00 / 11.64 | 80.00 / 29.49 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 80.00 / 39.03 | 40.00 / 74.73 | 44.44 / 87.91 | 88.89 / 15.63 | 80.00 / 20.32 | 80.00 / 18.73 |
@@
| FSDetView [Xiao2020FewShotOD] | - / - | 10.00 / 96.30 | - / - | - / - | - / - | - / - |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 55.56 / 35.79 | 55.00 / 39.50 | 25.00 / 59.75 | 33.33 / 61.73 | 10.00 / 75.44 | 50.00 / 29.89 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 77.78 / 19.83 | 90.00 / 15.53 | 50.00 / 34.62 | 50.00 / 40.90 | 50.00 / 42.89 | 100.00 / 17.54 |
@@
| FSDetView [Xiao2020FewShotOD] | - / - | - / - | - / - | - / - | - / - | - / - |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 50.00 / 32.32 | 22.22 / 53.95 | 30.00 / 86.84 | 20.00 / 69.35 | 50.00 / 45.98 | 20.00 / 59.59 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 70.00 / 22.11 | 88.89 / 16.71 | 30.00 / 50.37 | 50.00 / 48.06 | 70.00 / 28.76 | 40.00 / 47.36 |
@@
| FSDetView [Xiao2020FewShotOD] | - / - | - / - | - / - | 0.00 / 110.48 | - / - | 28.57 / 49.01 |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 77.78 / 31.87 | 50.00 / 48.89 | 0.00 / 51.62 | 30.00 / 67.83 | 0.00 / 101.22 | 14.29 / 53.71 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 66.67 / 22.15 | 90.00 / 21.53 | 80.00 / 24.69 | 50.00 / 35.85 | 50.00 / 74.92 | 14.29 / 54.06 |
@@
| FSDetView [Xiao2020FewShotOD] | - / - | 62.50 / 42.29 | - / - | - / - | - / - | - / - |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 20.00 / 73.80 | 75.00 / 19.83 | 50.00 / 36.90 | 10.00 / 56.94 | 90.00 / 19.91 | 25.00 / 101.14 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 80.00 / 36.15 | 50.00 / 70.09 | 70.00 / 28.20 | 40.00 / 38.71 | 70.00 / 21.11 | 50.00 / 51.71 |
@@
| FSDetView [Xiao2020FewShotOD] | 10.00 / 80.55 | - / - | - / - | - / - | 33.33 / 102.13 | - / - |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 20.00 / 43.97 | 57.14 / 30.07 | 0.00 / 116.94 | 50.00 / 49.04 | 50.00 / 64.67 | 30.00 / 50.62 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 40.00 / 29.71 | 71.43 / 14.63 | 0.00 / 89.58 | 56.25 / 33.90 | 50.00 / 61.28 | 80.00 / 30.68 |
@@
| FSDetView [Xiao2020FewShotOD] | - / - | - / - | - / - | - / - | 40.00 / 98.08 | 15.00 / 107.20 |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 50.00 / 66.47 | 62.50 / 63.46 | 20.00 / 50.26 | 75.00 / 34.54 | 50.00 / 62.30 | 75.00 / 33.12 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 50.00 / 23.66 | 87.50 / 32.48 | 40.00 / 45.30 | 100.00 / 16.06 | 50.00 / 32.59 | 65.00 / 39.84 |
@@
| FSDetView [Xiao2020FewShotOD] | - / - | 0.00 / 93.15 | - / - | - / - | - / - | |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 75.00 / 32.67 | 0.00 / 117.58 | 80.00 / 23.95 | 0.00 / 95.34 | 44.44 / 47.62 | |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 75.00 / 37.13 | 50.00 / 70.34 | 100.00 / 11.94 | 30.00 / 45.89 | 44.44 / 39.14 | |
@@
| FSDetView [Xiao2020FewShotOD] | 9.09 / 90.67 | 4.76 / 81.40 | 0.00 / 81.72 | 54.84 / 40.24 | 0.00 / 93.31 |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 4.55 / 83.23 | 0.00 / 78.44 | 6.67 / 74.79 | 3.23 / 87.56 | 3.03 / 74.11 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 9.09 / 78.51 | 0.00 / 78.95 | 13.33 / 76.79 | 6.45 / 81.73 | 3.03 / 85.85 |
@@
| FSDetView [Xiao2020FewShotOD] | 0.00 / 96.13 | 0.00 / 103.08 | |||
| Orient Anything [Wang2024OrientAL] (Vit-S) | 0.00 / 90.53 | 2.63 / 76.98 | |||
| Orient Anything [Wang2024OrientAL] (Vit-L) | 17.24 / 70.88 | 15.79 / 68.49 | |||
@@ B.4 Impact of Occlusion
As illustrated in Figure 24, our object generation module is capable of producing reasonable results under partial occlusion, with minimal performance degradation. However, in scenarios involving severe occlusion, we recommend using the method proposed in [Wu2025Amodal3RA3], which is specifically designed for occlusion-robustness, as the pre-trained checkpoint prior to applying our orientation-alignment fine-tuning. It is also important to note that our orientation estimation approach has not been designed to handle occluded inputs. Further research is necessary to extend its applicability to occlusion-prone scenarios.
@@ Appendix C License and Border Impact
| Objaverse [Deitke2022ObjaverseAU] | ODC-By v1.0 license | texthttps://objaverse.allenai.org/ |
| Trellis [Xiang2024Structured3L] | MIT License | texthttps://github.com/microsoft/TRELLIS |
| Wonder3D [Long2023Wonder3DSI] | MIT License | texthttps://github.com/xxlong0/Wonder3D/tree/main |
| FoundationPose [Wen2023FoundationPoseU6] | NVIDIA Source Code License | texthttps://github.com/NVlabs/FoundationPose |
| LGM [Tang2024LGMLM] | MIT License | texthttps://github.com/3DTopia/LGM |